# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v6

@@ 9 Related Work
This is the first paper to enable DRAM chips that autonomously and efficiently perform various maintenance operations with simple changes to existing DRAM interfaces. No prior work proposes setting the MC free from managing DRAM maintenance operations nor studies the system-level performance and energy impact of autonomous maintenance mechanisms. We briefly discuss relevant prior works.
Changing the DRAM Interface. Several prior works [129, 130, 131, 132] propose using high-speed serial links and packet-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose five new SMD-based maintenance mechanisms (section 5). Second, prior works significantly overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overhead. Many previous works [16, 17, 15, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] propose techniques to reduce DRAM refresh overhead. SMD enables maintenance mechanisms for reducing DRAM refresh overhead and other efficient maintenance mechanisms for improving DRAM reliability and security (section 5).
RowHammer Protection. Many prior works [133, 49, 50, 54, 48, 53, 134, 45, 135, 10, 75, 136, 59, 58, 57, 137, 62, 63, 138, 139] propose techniques for RowHammer protection. One can use SMD to implement existing or new RowHammer protection mechanisms. Our goal is not to prove that SMD-DRP is the best RowHammer mitigation mechanism.
Memory Scrubbing. Although prior works report that the overhead of memory scrubbing is small as low scrubbing rate (e.g., scrubbing period of 24 hours [69, 70, 72], 45 minutes per 1 GB scrubbing rate [66], only when idle [67]) is generally sufficient, the cost of scrubbing can dramatically increase for future DRAM chips due to increasing DRAM bit error rate and increasing DRAM chip density. SMD-MS enables efficient scrubbing by eliminating off-chip data transfers.
Leveraging Subarray-level Parallelism. Prior works overlap the latency of accessing multiple subarrays by modifying the DRAM architecture [79, 80, 16]. SMD’s maintenance-access parallelization builds on the basic design proposed in SALP [79] and refresh-access parallelization introduced in [16, 80]. A DDRx interface based on these prior works [79, 80, 16] still needs to change for each new maintenance operation that the DRAM manufacturer needs to implement. SMD, in contrast, allows the implementation of new maintenance mechanisms without making any further changes to the DDRx interface.
@@ 10 Conclusion
To set the memory controller free from managing DRAM maintenance operations, Self-Managing DRAM (SMD) introduces minimal changes to the existing DRAM chips and memory controllers. SMD enables in-DRAM maintenance operations with no further changes to the DRAM interface, memory controller, or other system components. Using SMD, we implement efficient maintenance mechanisms for DRAM refresh, RowHammer protection, and memory scrubbing. We show that these mechanisms altogether enable a higher performance, more energy efficient and at the same time more reliable and secure DRAM system. We believe and hope that SMD will enable practical adoption of innovative ideas in DRAM design.

### v7

@@ 10 Related Work
This is the first work that gives DRAM chips the ability (i.e., the breathing room) to autonomously perform maintenance operations with low-cost simple changes to the existing rigid DRAM interface. No prior work proposes setting the MC free from managing DRAM maintenance operations nor studies the system-level performance and energy impact of autonomous DRAM maintenance mechanisms. We briefly discuss relevant prior works.
Changing the DRAM Interface. Several prior works [218, 219, 220, 221, 222] propose using high-speed serial links and packet-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose eight SMD-based maintenance mechanisms (section 5) and rigorously evaluate three (section 5.1, section 5.2, and section 5.3) of them. Second, prior works significantly modify or even overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overheads. Many previous works [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] propose techniques to reduce performance and energy overheads of DRAM refresh. SMD can seamlessly integrate many of these techniques along with other efficient maintenance operations (section 5).
Read Disturbance Protection. Many prior works [47, 48, 49, 50, 12, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 18, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122] propose techniques for RowHammer and RowPress [159] protection. One can use SMD to implement these or new RowHammer protection mechanisms.
Memory Scrubbing. Various prior works report that the overhead of memory scrubbing is small as low scrubbing rates (e.g., scrubbing period of 24 hours [128, 129, 119], 45 minutes per 1 GB scrubbing rate [125], only when idle [126]) are generally sufficient. However, the cost of scrubbing can increase for future DRAM chips due to increasing DRAM bit error rate [179, 5, 8] and increasing DRAM chip density [17]. SMD-MS enables efficient scrubbing by eliminating off-chip data transfers.
Leveraging Subarray-level Parallelism. Prior works overlap the latency of accessing multiple subarrays by modifying the DRAM architecture [148, 149, 19]. SMD’s maintenance-access parallelization builds on the basic design proposed in SALP [148] and refresh-access parallelization introduced in [19, 149]. A DRAM interface based on these prior works [148, 149, 19] still needs to be modified for each new maintenance operation that the JEDEC standard dictates DRAM to implement. SMD, in contrast, allows the implementation of new maintenance mechanisms with a single, simple change to the DRAM interface.
@@ 11 Conclusion
We introduced a new, low-cost framework, Self-Managing DRAM (SMD), for accelerating the adoption of new in-DRAM maintenance operations. With a one-time low-cost modification to the DRAM interface, SMD enables implementing new in-DRAM maintenance operations with no further changes to the DRAM interface, memory controller, or other system components. Using SMD, we implement efficient in-DRAM maintenance mechanisms for DRAM refresh, RowHammer protection, and memory scrubbing. We show that these mechanisms enable a higher performance, more energy efficient, and, at the same time, more robust DRAM system. We believe and hope that SMD can enable practical adoption of future innovative ideas in DRAM design and inspire better ways of partitioning work between memory and processor chips.
Acknowledgments
We thank the anonymous reviewers of MICRO 2022, HPCA 2023, ISCA 2023, MICRO 2023, HPCA 2024, ISCA 2024, and MICRO 2024 for the feedback. We thank the SAFARI Research Group members for their valuable and constructive feedback along with the stimulating scientific and intellectual environment they provide. We acknowledge the generous gift funding provided by our industrial partners (especially Google, Huawei, Intel, Microsoft, VMware), which has been instrumental in enabling the research we have been conducting on memory systems. This work was also in part supported by the Google Security and Privacy Research Award, and the Microsoft Swiss Joint Research Center.